a charge of 20 x 10-6 c is 20 cm from another charge of 60 x 10-6 c. (b) determine the ep midway between them.

1
Expert's answer
2022-02-24T09:04:44-0500

"E=\frac{4k(q_1-q_2)}{r^2}=36~\frac{MV}m."